there are two more questions which i dint get.....the questions are as follows:-


1]the pressure of  liquid column depends on its height h, the density p[its the symbol of density],acceleration due to gravity g,show that p=h [symbol of density]p g by dimensional method.(give constant k=1)


2]express a pressure of 50 N/metre square in terms of dyne/cm square.


urgently needed for the test tht is conducted tomorrow...............

Grade:Upto college level

3 Answers

AskiitianExpert Shine
14 years ago

Hi

Take, P= k x ha x Þb x gc .

dimensions of P are kg/ms2 . Write the dimensions of h, g & Þ .

h=m , g=m/s2 , Þ= kg/m3

Equate the power of each i.e of m, kg and sec on the left hand side to rt hand side. U ll get three equations, solve them for three variables. U ll get the final answer as a=b=c=1 .

therefore P=hÞg as k=1

AskiitianExpert Shine
14 years ago

Ans 2:

Units of force
  newton
(SI unit)
dyne kilogram-force,
kilopond
pound-force poundal
1 N ≡ 1 kg·m/s² = 105 dyn ≈ 0.10197 kp ≈ 0.22481 lbf ≈ 7.2330 pdl
1 dyn = 10−5 N ≡ 1 g·cm/s² ≈ 1.0197×10−6 kp ≈ 2.2481×10−6 lbf ≈ 7.2330×10−5 pdl
1 kp = 9.80665 N = 980665 dyn gn·(1 kg) ≈ 2.2046 lbf ≈ 70.932 pdl
1 lbf ≈ 4.448222 N ≈ 444822 dyn ≈ 0.45359 kp gn·(1 lb) ≈ 32.174 pdl
1 pdl ≈ 0.138255 N ≈ 13825 dyn ≈ 0.014098 kp ≈ 0.031081 lbf ≡ 1 lb·ft/s²

1 Newton=  105dynes and 1 m = 100cm , therefore 1m2 = 104cm2 .

therefore 50N/m sq = 50 x 105 / 104 dynes /cm sq, hence answer is 500.

anshul dixit
13 years ago

u can use..........M1/M2=1000  L1/L2=100 T1/T2=1

n1ui=n2u2..........take the dymensional formula of newton/metresquare...